OpenStudy (anonymous):

What percent of 159 is 129?

OpenStudy (anonymous):

Let the percentage be x.. \[x \% \; of \; \; 159 = 129\]

OpenStudy (jiteshmeghwal9):

Let the percent be x% then x% of 159=129 \(\frac{x}{100}\times 159=129\)

OpenStudy (anonymous):

\[\frac{x}{100} \times 159 = 129\] \[x = \frac{129 \times 100}{159} = Solve \; \; this..\]

OpenStudy (jiteshmeghwal9):

now solve . I think i have made this much easier . From here @shivraj can easily do it :)

OpenStudy (anonymous):

Your final answer results in 18.86 which is incorrect.. @heedcom

OpenStudy (anonymous):

81

OpenStudy (anonymous):

\[\frac{18.86}{100} \times 159 = 29.98 = Incorrect\]

OpenStudy (anonymous):

You should get 129 and not 29.98 @heedcom

OpenStudy (jiteshmeghwal9):

@waterineyes is correct because i checked this in calculator also :)

OpenStudy (anonymous):

ok, cool, sorry guys, i did it the other way round.

OpenStudy (anonymous):

Precisely my method is correct @jiteshmeghwal9 And yours too. He he he..

OpenStudy (jiteshmeghwal9):

np :)

OpenStudy (jiteshmeghwal9):

LOL ;)

OpenStudy (anonymous):

Are you clear with this @shivraj

OpenStudy (anonymous):

yes thnks